OpenStudy (anonymous):

V = C^2[0,1] is defined as the space of twice-differentiable real-valued functions on [0,1]. I'm wondering what it means to be on [0,1]. I think it has to do with the inner product. (i.e. the bounds of integration for the inner product are 0,1). What would be the most general function from V?

OpenStudy (mathteacher1729):

"\(V = \mathcal{C}^2 \ [0,1]\) is defined as the space of twice-differentiable real-valued functions on \([0,1]\)." means V is the set of all functions which are twice-differentiable on the closed interval \(0\leq x \leq 1\). That means that if you have some function \(f\) on this interval, then \(f' \ \text{ and } \ f''\) exist and are well-defined on this interval as well. Some examples: f(x) = x f(x) = x^2 f(x) = (any polynomial) f(x) = sin(x) f(x) = cos(x) f(x) = e^x

OpenStudy (anonymous):

Thanks for the response. That makes sense, but the reason I asked for the most general function is that I am asked to compute two linearly independent eigenvectors for a genral eigenvalue x > 0.

OpenStudy (mathteacher1729):

eigenvectors in this case are functions (the FUNCTIONS THEMSELVES are considered vectors in this vector space...) which are not constant multiplies of one another. Usually \(f(x)=c_1\sin(x)+c_2\cos(x)\) (the function f is come linear combination of two linearly independent functions sin and cosine) is a classic example of this. \(f(x) = c_1x+c_2x^2\) is another one.

OpenStudy (mathteacher1729):

Also \[\Huge f(x) = c_1e^{mx}+c_2e^{nx}\] where \(m\neq n\)

OpenStudy (anonymous):

This is very clear now. Thank you.

OpenStudy (mathteacher1729):

Question -- have you studied linear algebra?

OpenStudy (anonymous):

I have taken an introductory linear algebra course, and I am now in an applied linear course. I'm used to calculating eigenvalues/eigenvectors for matrices and transformations that have explicit matrix representations, however this question requires a deeper understanding.

OpenStudy (mathteacher1729):

Check out Ch 02 (Vector Spaces) from Jim Hefferon's FREE text (with lots of examples, and SOLVED problems) http://joshua.smcvt.edu/linalg.html/ it might be helpful. :)

OpenStudy (anonymous):

Wow this looks great. It will make a good complement to the Linear Algebra Schaum's outline that we use for the course.